Home has 1200 units of labor available. It can produce two goods as Apples and Bananas.

The unit labour requirement in apple production is 3. For Banana production ,it is 2.

1. Graph Home's Production Possibility Frontier

2. What is opportunity cost for apples in terms of bananas?

3. In the absence of trade, what would the price of apples in terms of bananas? Why?


1
Expert's answer
2022-02-07T17:02:12-0500

Home has 1200 units of labor available. It can produce two goods as Apples and Bananas.

The unit labor requirement in apple production is 3. For Banana production, it is 2.

1. Graph Home's Production Possibility Frontier

Home can produce apples = "\frac{1200}{3}" = 400 apples

Home can produce bananas = "\frac{1200}{2}" = 600 bananas

Therefore, the Home's Production Possibility Frontie is:




2. What is opportunity cost for apples in terms of bananas?

The opportunity cost of producing apples in home = "\frac{600}{400}" =1.5,

Therefore, the opportunity cost of producing apples in terms of bananas is 1.5.


3. In the absence of trade, what would the price of apples be in terms of bananas? Why?

The price of apples be in terms of bananas would be 1.5 because, at equilibrium, cost should be equal to the price to achieve efficiency.